Jason has 4 more dimes than nickels and 2 fwere than 5 times as many quarters as nickels. He has $9.70. How many dimes does he have.
Answer by jorel1380(3719)

Let n=nickels
dimes=n+4
quarters=5n-2
5n+10(n+4)+25(5n-2)=970
5n+10n+40+125n-50=970
140n=980
n=7
d=11
q=33
Jason has 7 nickels, 11 dimes, and 33 quarters..